# Weinstein exactness of nearby Lagrangians and the Lagrangian $C^{0}$ flux conjecture

### v2

@@ 4.2 Finiteness of the capacities
Let $L$ be a Lagrangian submanifold of $M$. Suppose that, as an abstract manifold, $L$ admits a Lagrangian embedding in a Liouville domain $W$ with $SH(W)=0$. For every $\tau\geq 0$, there exists a Weinstein neighbourhood $\mathcal{W}(L)$ of $L$ in $M$, such that if $L^{\prime}\in\mathcal{L}(\tau)$ is included in $\mathcal{W}(L)$, then the map
${H_{1}(L^{\prime};\mathds{R})}$${H_{1}(L;\mathds{R})}$$\scriptstyle{\pi_{*}}$
induced by the projection $\pi:L^{\prime}\to L$ is nonzero.
In turn, this follows from a variant of Proposition 33 and a proof of the finiteness of $c^{0}_{\mathrm{all}}(D^{*}_{R}L)$ for $L$ as in the theorem. More precisely, we need the following two results.
@@
Let $L$ be a Lagrangian in a symplectic manifold $M$. Suppose that $c_{Q}^{0}(D^{*}_{R}L)$ is finite for some $R>0$. For every $\tau\geq 0$, there exists a Weinstein neighbourhood $\mathcal{W}(L)$ of $L$ in $M$, such that all $L^{\prime}\in\mathcal{L}(Q,\tau)$ included in $\mathcal{W}(L)$ has nontrivial morphism $\pi_{*}:H_{1}(L^{\prime};\mathds{R})\to H_{1}(L;\mathds{R})$ if $H_{1}(L;\mathds{R})\neq 0$.
If $c_{\mathrm{all}}^{0}(D^{*}_{R}L)$ is finite for some $R>0$, then the same holds for all $L^{\prime}\in\mathcal{L}(\tau)$.
@@
The proof is essentially that of Proposition 33, except that Observation (2) is replaced with the following.
  • (2’)
    If $L^{\prime}\in\mathcal{L}(Q,\tau)$, $L^{\prime}\subseteq\mathcal{W}_{r}(L)$, and $H_{1}(L^{\prime};\mathds{R})\to H_{1}(\mathcal{W}(L);\mathds{R})$ is zero, then $c_{Q}^{0}(D^{*}_{r}L)\geq\tau$ whenever we have that $\omega(H_{2}(\mathcal{W}_{r}(L),L^{\prime}))\neq 0$.
Therefore, for $r$ small enough, all $L^{\prime}\in\mathcal{L}(Q,\tau)$ such that $L^{\prime}\subseteq\mathcal{W}_{r}(L)$ must be either $H$-exact or induce a nonzero map $H_{1}(L^{\prime};\mathds{R})\to H_{1}(\mathcal{W}_{r}(L);\mathds{R})$.
By the central lemma 24, $H$-exactness in $\mathcal{W}_{r}(L)$ implies that the projection $\pi:L^{\prime}\to L$ is a homotopy equivalence. Therefore, it also implies that the map $H_{1}(L^{\prime};\mathds{R})\to H_{1}(\mathcal{W}_{r}(L);\mathds{R})$ is nonzero whenever $H_{1}(\mathcal{W}_{r}(L);\mathds{R})\neq 0$. Since the projection $\mathcal{W}_{r}(L)\to L$ is a homotopy equivalence, this implies the result.
Lemma 38
Let $L$ be a manifold which admits a Lagrangian embedding in $W$ with $SH(W)=0$. Then,
$c_{\mathrm{all}}^{0}(D^{*}_{R}L)<\infty$
for some $R>0$.
Proof.
By the Weinstein neighbourhood theorem, there is some $R>0$ such that $D^{*}_{R}L$ embeds symplectically in $W$. Therefore, we have that
$c_{\mathrm{all}}^{0}(D^{*}_{R}L)\leq Bc^{0}_{\mathrm{all}}(W)<\infty$
where $B=B(L)$ is the constant of Lemma 30. The latter finiteness is that of Theorem 35 above.
We now get pretty directly proofs of Theorems 36 and 3, Case (a).
@@
Combining Proposition 37 and Lemma 38 gives the existence of a neighbourhood $\mathcal{W}(L)$ such that whenever $L^{\prime}\in\mathcal{L}(\tau)$ is in $\mathcal{W}(L)$, then $H_{1}(L^{\prime};\mathds{R})\to H_{1}(L;\mathds{R})$ is nonzero if $H_{1}(L;\mathds{R})\neq 0$. But by the Viterbo transfer morphism, $H_{1}(L;\mathds{R})\neq 0$ since it embeds as a Lagrangian in a Liouville domain with $SH(W)=0$ (see [Rit13]). ∎
Proof of Theorem 3, Case (a).
Whenever $\dim H_{1}(Q;\mathds{R})=1$ and $L$ admits a Lagrangian embedding in $W$ with $SH(W)=0$, then
$c_{Q}(D^{*}_{R}L)=c^{0}_{Q}(D^{*}_{R}L)\leq c^{0}_{\mathrm{all}}(D^{*}_{R}L)<\infty,$
where the first equality is Lemma 31 and the last inequality is Lemma 38. Case (a) of Theorem 3 then follows from Proposition 33. ∎
Remarks 8.
In view of Lemma 32, it actually suffices to prove finiteness of $c^{0}$ for $Q\times Q^{\prime}$ admitting an embedding in $W$ with $SH(W)=0$, where $H_{1}(Q^{\prime};\mathds{R})=0$. However, we cannot find an example of a $Q$ with $H_{1}(Q;\mathds{R})=\mathds{R}$ such that $Q\times Q^{\prime}$ admits such an embedding but not $Q$ by itself.
4.3 Lagrangian Klein bottles in cotangent bundles
We now focus our efforts on proving Case (c) of Theorem 3, i.e. the case of the Klein bottle $K$. The proof is fairly different from previous cases, as it relies on the classification of Lagrangian Klein bottles in $T^{*}K$, which turns out to be doable by a direct computation. It does however rely on the deep fact that there is no Lagrangian Klein bottle in $\mathds{C}^{2}$ [She09].
Theorem 39
Proof.
Using the flat metric on $\mathds{T}^{2}$ and $K$, the 2:1 covering $p:\mathds{T}^{2}\to K$ lifts to another 2:1 covering $\widetilde{p}:T^{*}\mathds{T}^{2}\to T^{*}K$ which is also a local isometry and symplectomorphism. Therefore, $\widetilde{L}:=\widetilde{p}^{-1}(L)$ must be a (possibly disconnected) Lagrangian submanifold of $D^{*}_{r}\mathds{T}^{2}$. Since $\widetilde{p}|_{\widetilde{L}}$ is also a 2:1 covering, $\widetilde{L}$ must either be two disconnected copies of a Klein bottle or a 2-torus. However, if the former was the case, then each connected component of $\Psi(\widetilde{L})$ would be a Lagrangian Klein bottle in $\mathds{C}^{2}$, which does not exist [She09]. Therefore, $\widetilde{L}$ must be a 2-torus. In other words, the composition
[…]
admits a lift to $T^{*}\mathds{T}^{2}$, but the composition
[…]
does not.
5 Theorem 4 and the $C^{0}$ Lagrangian flux conjecture
In this section, we first prove Theorem 4 (Section 5.1). We then give a short proof of Theorem 6 (Section 5.2), which follows almost directly from the proof of Theorem 4. Finally, we prove a refined version of Proposition 7 and use it to properly show Corollary 8 (Section 5.3).
@@ 5.1.2 Proof of Lemma 40
In case (b), note that $H_{1}(L;\mathds{R})\to H_{1}(M;\mathds{R})$ being zero is equivalent to $H_{1}(L)\to H_{1}(M)$ being finite, since $H_{1}(\,\cdot\,;\mathds{R})=H_{1}(\,\cdot\,)\otimes\mathds{R}$. By the long exact sequence of the pair $(M,L)$, this is in turn equivalent to $H_{2}(M,L)\to H_{1}(L)$ having finite cokernel, whose size we denote by $k$. Then, $k\beta$ bounds some $u\in H_{2}(M,L)$, and we have that
@@ 5.2 Proof of Theorem 6
We now turn our attention to Theorem 6 on limits of $H$-rational Lagrangians. As we shall see, the theorem follows pretty directly from the techniques that we developed to prove Theorem 3 and 4.
@@
The second part — that is, for when we know that the $L_{i}$’s are $H$-exact in $\mathcal{W}(L)$ for $i$ large — follows from the proof of Theorem 4. Indeed, the proof of Lemma 40 gives that $\lambda_{0}|_{L_{i}}$ takes values in $\tau_{i}\mathds{Z}$ on the image of the boundary morphism $\partial_{i}:H_{2}(M,L_{i})\to H_{1}(L_{i})$. Here, we identify $L_{i}$ with its preimage in $T^{*}L$ under a Weinstein neighbourhood of $L$. Just like in Theorem 4, we can thus use Proposition 41 to conclude that $\lambda_{0}(\partial_{i}(H_{2}(M,L_{i})))=0$ if $L_{i}\subseteq\mathcal{W}_{r}(L)$ for $r$ small enough, i.e. for $i$ large enough. Note that $r$ may be taken independently of $i$ since $\inf\tau_{i}>0$. But then, this means that, for all $A\in H_{2}(M,L_{i})$,
$\omega(A)=\omega(A\#C)-\omega(C)=\omega(A\#C)-\lambda_{0}(\partial C)=\omega(A\#C)\in\omega(H_{2}(M,L)),$
where $C$ is the usual (union of) cylinder in $T^{*}L$ from $\partial A$ to $\pi(\partial A)$ and $\pi:T^{*}L\to L$ the canonical projection. Therefore, we have that $\omega(H_{2}(M,L_{i}))\subseteq\omega(H_{2}(M,L))$. But $H$-exactness implies that $L_{i}\to T^{*}L\to L$ is a homotopy equivalence by Lemma 24. Therefore, for every $A\in H_{2}(M,L)$, $\partial A$ admits a lift $a\in H_{1}(L_{i})$. We can then again create a cylinder $C$ from $\partial A$ to $a$ and run the above argument to get $\omega(H_{2}(M,L))\subseteq\omega(H_{2}(M,L_{i}))$ for $i$ large. Therefore, we have that
$\omega(H_{2}(M,L))=\omega(H_{2}(M,L_{i}))=\tau_{i}\mathds{Z}$
for large $i$. This is only possible if $\tau_{i}$ is independent of $i$ for $i$ large. ∎
5.3 Proof of Proposition 7

### v3

@@ 4.3 Proof of Theorem 4
Let $L$ be a Lagrangian submanifold in $M$. Suppose that, as an abstract manifold, $L$ admits a Lagrangian embedding in a Liouville domain $W$ with $SH(W)=0$. For every $\tau\geq 0$, there exists a Weinstein neighbourhood $\mathcal{W}(L)$ of $L$ in $M$, such that if $L^{\prime}\in\mathcal{L}(\tau)$ is included in $\mathcal{W}(L)$, then the map $\pi_{*}:H_{1}(L^{\prime};\mathds{R})\to H_{1}(L;\mathds{R})$ induced by the projection is nonzero.
Sketch of proof.
By adapting the proof of Proposition 32, one can see that finiteness of $c_{\mathrm{all}}^{0}(D^{*}_{R}L)$ for some $R>0$ is equivalent to the following version of Conjecture C: if $K$ is an $H$-rational Lagrangian inside $D^{*}_{r}L$ such that the map $(\pi|_{K})_{*}:H_{1}(K;\mathds{R})\to H_{1}(L;\mathds{R})$ is zero, then the $H$-rationality constant $\tau_{K}$ of $K$ satisfies $\tau_{K}\leq Cr$ for a constant $C=C(L)$.
However, the finiteness of $c_{\mathrm{all}}^{0}(D^{*}_{R}L)$ for some $R>0$ follows directly from the Weinstein neighbourhood theorem, the monotonicity of the $c^{0}$ capacities (Lemma 29), and Zhou’s Theorem.
Replicating the proof of Theorem 4, one gets the existence of a neighbourhood $\mathcal{W}(L)$ such that whenever $L^{\prime}\in\mathcal{L}(\tau)$ is in $\mathcal{W}(L)$, then $H_{1}(L^{\prime};\mathds{R})\to H_{1}(L;\mathds{R})$ is nonzero. ∎
Remark 11.
Since $L$ embeds as a Lagrangian in a Liouville domain with $SH(W)=0$, so does any nearby $L^{\prime}$. The Viterbo transfer morphism then implies that $H_{1}(L;\mathds{R})$ and $\ H_{1}(L^{\prime};\mathds{R})$ are nonzero (see [Rit13]). Therefore, $H_{1}(L^{\prime};\mathds{R})\to H_{1}(L;\mathds{R})$ being nonzero is consistent with the hypotheses on $L$ and $L^{\prime}$.
4.4 Conjecture C and covering spaces
The condition on the Lagrangian $K\subset D^{*}_{r}L$ in Conjecture C, i.e. non-surjectivity of $(\pi_{K})_{*}$, while technical, has the advantage of behaving well with covering spaces. This is highlighted by the following proposition, which proves the second part of Theorem 2.
@@ 4.4 Conjecture C and covering spaces
Suppose that Conjecture C holds for a closed connected manifold $L^{\prime}$, then it holds for all $L$ such that there exists a covering map $\pi:L^{\prime}\rightarrow L$.
Before we prove Proposition 37, we require the following lemma from group theory. For a group $G$, denote by ${\mathrm{Ab}}(G)=G/[G,G]$ its abelianization.
Lemma 38
Suppose $H$ is a subgroup of finite index in a group $G.$ Then the inclusion $i:H\to G$ induces a surjection
$\operatorname{{{\mathrm{Ab}}}}(H)\otimes\mathds{Q}\to\operatorname{{{\mathrm{Ab}}}}(G)\otimes\mathds{Q}.$
@@
Observe that $\operatorname{\mathrm{rank}}\operatorname{{{\mathrm{Ab}}}}(G)=\dim_{\mathds{Q}}\operatorname{Hom}(G,\mathds{Q}).$ It thus suffices to prove that the restriction map
$r:\operatorname{Hom}(G,\mathds{Q})\to\operatorname{Hom}(H,\mathds{Q})$
is injective. Suppose $\nu\in\ker(r),$ i.e. $\nu|_{H}\equiv e_{H}$. Let $H=a_{1}H,a_{2}H,\ldots,a_{k}H$ be the right cosets of $H$ in $G.$ Then $\operatorname{\mathrm{im}}(\nu)=\{e_{H},\nu(a_{2}),...,\nu(a_{k})\}\subset\mathds{Q}$ is a finite subset. Since $\mathds{Q}$ has no non-zero elements of finite order, this yields that $\nu\equiv e_{H}.$
Proof of Proposition 37.
First, extend the covering $p:L^{\prime}\to L$ to a covering
$p:W^{\prime}_{r}=D^{*}_{r}L^{\prime}\to W_{r}=D^{*}_{r}L$
[…]
It is now sufficient to prove that if $(\pi_{K^{\prime}})_{*}:H_{1}(K^{\prime};\mathds{Q})\to H_{1}(L^{\prime};\mathds{Q})$ is surjective, then $(\pi_{K})_{*}:H_{1}(K;\mathds{Q})\to H_{1}(L;\mathds{Q})$ also is. Note that as covering maps, $p$ and $p_{K^{\prime}}$, induce injective homomorphisms at the level of fundamental groups. As $H_{1}(X)=\operatorname{{{\mathrm{Ab}}}}\pi_{1}(X)$ for any path connected space $X$, Corollary 38 implies that $p_{*}:H_{1}(L^{\prime};\mathds{Q})\to H_{1}(L;\mathds{Q})$ and $(p_{K})_{*}:H_{1}(K^{\prime};\mathds{Q})\to H_{1}(K;\mathds{Q})$ are surjective. Hence if $(\pi_{K^{\prime}})_{*}:H_{1}(K^{\prime};\mathds{Q})\to H_{1}(L^{\prime};\mathds{Q})$ is surjective, then $(p\circ\pi_{K^{\prime}})_{*}=p_{*}\circ(\pi_{K^{\prime}})_{*}:H_{1}(K^{\prime};\mathds{Q})\to H_{1}(L^{\prime};\mathds{Q})$ is surjective. However, as $p\circ\pi_{K^{\prime}}=\pi_{K}\circ p_{K^{\prime}}$ the image of $(p\circ\pi_{K^{\prime}})_{*}=(\pi_{K}\circ p_{K^{\prime}})_{*}=(\pi_{K})_{*}\circ(p_{K^{\prime}})_{*}$ is contained in that of $(\pi_{K})_{*}.$ Therefore $(\pi_{K})_{*}$ is surjective. ∎
4.5 Homotopy version of Theorem 2
The argument in Proposition 37 implies that the homotopy version of Conjecture C is also invariant under coverings: if the rationality constant of all $K^{\prime}\subseteq D^{*}_{r}L^{\prime}$ with $(\pi_{K^{\prime}})_{*}:H_{1}(K^{\prime};\mathds{Q})\to H_{1}(L^{\prime};\mathds{Q})$ nonsurjective· is uniformly bounded and $L^{\prime}\to L$ is a covering, then the rationality constant of all $K\subseteq D^{*}_{r}L$ with $(\pi_{K})_{*}:H_{1}(K;\mathds{Q})\to H_{1}(L;\mathds{Q})$ nonsurjective is uniformly bounded by the same constant. As noted in Remark 8, this thus gives the homotopy version of Conjecture C for anything covered by $L_{0}\times\mathds{T}^{m}$ with $\pi_{1}(L_{0})$ finite.
Let us however finally give a direct proof of the homotopy version of Theorem 2, which we now state precisely.
Theorem 39
Suppose that $L$ is diffeomorphic to the product $L_{0}\times\mathds{T}^{m}$ of a closed manifold $L_{0}$ with $H_{1}(L_{0};\mathds{R})=0$ and a $m$-torus. Suppose that $K$ is a rational Lagrangian in $D^{*}_{r}L$ with rationality constant $\tau_{K}$ and such that $(\pi_{K})_{*}$ is not surjective, then
$\tau_{K}\leq Cr$
for a constant $C>0$ depending only on a choice of an auxiliary metric on $L$. Moreover, the same is true for any closed $L^{\prime}$ which is covered by $L$.
@@ 4.5 Homotopy version of Theorem 2
Suppose $f$ is a rational Lagrangian embedding of $K$ into
$D^{*}_{a}(T^{m}\times Q)\cong(D^{*}_{a}S^{1})^{m}\times D^{*}_{a}Q,$
with rationality constant $\tau_{K}>0$. Observe that, as an open symplectic manifold, $D^{*}_{a}S^{1}=(-a,a)\times S^{1}$ is symplectomorphic to the punctured disk $D^{2}(2a)^{\prime}=D^{2}(2a)\setminus\{0\}$ of area $2a$. The symplectomorphism is given by
$\psi:D^{*}_{a}S^{1}\to D^{2}(2a)^{\prime},\quad(p,q)\mapsto(r,\theta)$
where $\pi r^{2}=a+p$ and $\theta=2\pi q$ — this identifies the zero section $0_{S^{1}}=\{0\}\times S^{1}$ with the circle $S(a)=\partial D(a).$ In particular $(D^{*}_{a}S^{1})^{m}\times D^{*}_{a}Q$ is symplectomorphic via $\psi^{m}\times\operatorname{\mathrm{id}}$ to $V=(D^{2}(2a)^{\prime})^{m}\times D^{*}_{a}Q.$
Consider the induced map $f_{*}:H_{1}(K)\to H_{1}(V)$. If $R=\ker(f_{*})=0,$ then $K$ is $H$-exact by the central lemma (Lemma 24), and $\tau_{K}=0$; see also Remark 1. Suppose therefore that $R=\ker(f_{*})\neq 0.$ Let
$V_{i}=(D^{2}(2a)^{\prime})^{i-1}\times D^{2}(2a)\times(D^{2}(2a)^{\prime})^{m-i}\times D^{*}_{\varepsilon}Q$
for $1\leq i\leq m$. Let $g_{i}:V\to V_{i}$ denote the inclusion. We claim that $R=\ker(f_{*})=\ker((g_{i}\circ f)_{*})$ for at least one index $i\in\{1,\ldots,m\}.$ Indeed, as the image $I\subset\mathds{Z}^{m}$ of $f_{*}$ is a free abelian subgroup of rank at most $m-1,$ there exists at least one such $i$ with the projection $\pi_{i}:\mathds{Z}^{m}\to Z_{i}=\{k\in\mathds{Z}^{m}\;|\;k_{i}=0\}$ being injective on $I.$
However $(g_{i})_{*}:H_{1}(V)\to H_{1}(V_{i})$ is naturally identified with $\pi_{i}.$ This means that the period groups
${\mathcal{P}}_{V,K}={\left<[\omega],H^{D}_{2}(V,K;\mathds{Z})\right>},\,{\mathcal{P}}_{V_{i},K}={\left<[\omega],H^{D}_{2}(V_{i},K;\mathds{Z})\right>},$
coincide. In particular $K$ is rational in $V_{i}$ with the same rationality constant $\tau_{K}.$ For topological considerations, the same is true for $K$ inside
$\widehat{V}_{i}=(D^{2}(2a)^{\prime})^{i-1}\times\mathds{C}\times(D^{2}(2a)^{\prime})^{m-i}\times D^{*}_{\varepsilon}Q.$
Now, as $K\subset V_{i}$, the displacement energy of $K$ inside $\widehat{V}_{i}$ is at most $2a.$ By [Che98], this yields $\tau_{K}\leq 2a.$
Note that we directly get Proposition 3 for rational Lagrangians in $\mathds{R}^{n}\times[-1,1]^{n}$ by setting $a=1$ and considering by contradiction the embedded copy of the rational Lagrangian $K$ inside $\mathds{R}^{n}/\mathds{Z}^{n}\times[-1,1]^{n}.$ Note, however, that in this case the map $f_{*}$ from the proof above vanishes identically, whence we may choose any index $i$ to run the argument. This implies that one may replace $[-1,1]^{n}$ by $X$ as in Proposition 3.
5 Theorem 5 and the Lagrangian $C^{0}$ flux conjecture
We start this section by proving Theorem 5 (Section 5.1). We then turn to the Lagrangian $C^{0}$ flux conjecture and first give a short proof of Theorem 7 (Section 5.2), which follows almost directly from the proof of Theorem 5. Finally, we prove a refined version of Proposition 8 and use it to properly show Corollary 9, that is, the Lagrangian $C^{0}$ flux conjecture (Section 5.3). We conclude the section with a proof of Proposition 15 (Section 5.4).
@@ 5.1.2 Proof of Lemma 40
In case (b), note that $H_{1}(L;\mathds{R})\to H_{1}(M;\mathds{R})$ being zero is equivalent to the image of $H_{1}(L)\to H_{1}(M)$ being finite, since $H_{1}(\,\cdot\,;\mathds{R})=H_{1}(\,\cdot\,)\otimes\mathds{R}$. By the long exact sequence of the pair $(M,L)$, this is in turn equivalent to $H_{2}(M,L)\to H_{1}(L)$ having finite cokernel, whose size we denote by $k$. Then, $k\beta$ bounds some $u\in H_{2}(M,L)$, and we have that
@@ 5.2 Proof of Theorem 7
We now turn our attention to Theorem 7 on limits of $H$-rational Lagrangians. As we shall see, the theorem follows pretty directly from the techniques that we developed to prove Theorem 4 and 5.
@@
The second part — that is, for when we know that the $L_{i}$’s are $H$-exact in $\mathcal{W}(L)$ for $i$ large — is proved in a very similar way as Theorem 5. Indeed, we know that the projection $L_{i}\to L$ is a homotopy equivalence by Lemma 24. In particular, any homology class $A\in H_{2}(M,L;\mathds{Z})$ can be obtained by gluing a class $A_{i}\in H_{2}(M,L_{i};\mathds{Z})$ to a (union of) cylinder $C_{i}$ with $\partial C=\partial A_{i}-\partial A$ and $\partial A_{i}=(\pi|_{L_{i}})^{-1}(\partial A)$. Here, we identify $L_{i}$ with its preimage in $T^{*}L$ under a Weinstein neighbourhood of $L$. If $L_{i}\subseteq D^{*}_{r_{i}}L$, then Proposition 41 gives
$\lim|\lambda_{0}(\partial A_{i})|\leq\lim r_{i}\ell^{\min}_{g}(\partial A)=0,$
where we have made use of the fact that we may take $\lim r_{i}=0$ since $\{L_{i}\}$ Hausdorff-converges to $L$. Therefore,
$|\omega(A)|=\liminf|\omega(A_{i})+\lambda_{0}(\partial A_{i})|=\liminf|\omega(A_{i})|.$
But by rationality, $|\omega(A_{i})|=n_{i}\tau_{i}$ for some $n_{i}\in\mathds{Z}_{\geq 0}$. Since $\inf\tau_{i}>0$, $\{n_{i}\}$ must be bounded. Therefore, by passing to a subsequence if necessary, we may suppose that $n_{i}\equiv n$ for all $i$. Thus, if $\tau:=\liminf\tau_{i}$, then $\omega(A)\in\tau\mathds{Z}$. In particular, $L$ is $H$-rational and its rationality constant $\tau_{L}$ is a multiple of $\tau$.
We now prove that this multiple must be in fact 1. Pick a base $\{A^{1},\dots,A^{k}\}$ of the free part of $H_{2}(M,L;\mathds{Z})$ such that $\omega(A^{j})=\tau_{L}$ for all $j$, and construct $A^{j}_{i}$ and $C^{j}_{i}$ as above. By the same logic as above, we may suppose that $|\omega(A_{i}^{j})|=n^{j}\tau_{i}$ for all $i$. Therefore, we get that
$n^{j}\tau_{i}-r_{i}\ell\leq\tau_{L}\leq n^{j}\tau_{i}+r_{i}\ell,$
where $\ell:=\max_{j}\ell_{g}^{\min}(\partial A^{j})$. Since $\inf\tau_{i}>0$, for $i$ large enough, $r_{i}\ell<\tau$, and thus we must have $\tau_{L}=n^{j}\tau$ for all $j$, i.e. $n\equiv n^{j}$. But since the projection $L_{i}\to L$ is a homotopy equivalence, $\{A^{1}_{i},\dots,A^{k}_{i}\}$ is a basis of the free part of $H_{2}(M,L_{i};\mathds{Z})$. This is only possible if $n=1$ as $|\omega(A_{i}^{j})|=n\tau_{i}$ and $\omega(H_{2}(M,L_{i};\mathds{Z}))=\tau_{i}\mathds{Z}$. This thus implies that $\tau_{L}=\tau$.
Finally, to conclude that $\lim\tau_{i}$ must exists, note that the contrary would imply the existence of two subsequences of $\{L_{i}\}$ — thus still converging to $L$ — such that their corresponding subsequences of $\{\tau_{i}\}$ converge to different values. But then, both these values would need to be equal to $\tau_{L}$, which is not possible. ∎
5.3 Proof of Proposition 8 and Corollary 9